OpenStudy (anonymous):

factor gcf 36x^ 3 - 30x ^2

OpenStudy (anonymous):

er no 6x2 because what would be left is 6x and for 30 it would be 5 left

OpenStudy (anonymous):

36=3*2*3*2*x*x*x 30=3*2*5*x*x \[GCF=3*2*x*x=6x^2\]